### v1

# Scaling-Up the Pretraining of the Earth Observation Foundation Model PhilEO to the MajorTOM Dataset

Abstract
Today, Earth Observation (EO) satellites generate massive volumes of data, with the Copernicus Sentinel-2 constellation alone producing approximately 1.6TB per day. To fully exploit this information, it is essential to pretrain EO Foundation Models (FMs) on large unlabeled datasets, enabling efficient fine-tuning for downstream tasks with minimal labeled data. In this work, we present the scaling-up of our recently proposed EO Foundation Model, PhilEO Geo-Aware U-Net, on the unlabeled 23TB dataset MajorTOM, which covers the vast majority of the Earth’s surface, as well as on the specialized subset FastTOM 2TB that does not include oceans and ice. We develop and study various PhilEO models with different numbers of parameters and architectures. Finally, we fine-tune the models on the PhilEO Bench for road density estimation, building density pixel-wise regression, and land cover semantic segmentation, and we evaluate their performance. Our results demonstrate that for all n-shots for road density regression, the PhilEO 44M MajorTOM 23TB model outperforms PhilEO Globe 0.5TB 44M. We also show that for most n-shots for road density estimation and building density regression, PhilEO 200M FastTOM outperforms all the other models. The effectiveness of both dataset and model scaling is validated using the PhilEO Bench. We also study the impact of architecture scaling, transitioning from U-Net Convolutional Neural Networks (CNN) to Vision Transformers (ViT).
Index Terms:
Earth Observation, Foundation Models.
I Introduction
Given the vast amounts of satellite data coming from multiple constellations, such as Copernicus Sentinel-2, the focus has recently been shifted to Self-Supervised Learning (SSL). Masked Autoencoders [1] and Contrastive Learning [2] aim at reducing the need for large labeled datasets and have led to a growing interest in Foundation Models (FM) for Earth Observation (EO) [4], [5]. FMs are pretrained on satellite data to learn robust feature representations and, then, are fine-tuned on smaller labeled datasets for various downstream tasks. FMs use different architectures for the encoding (and decoding) of the data and features, which lead to different performance results for the downstream tasks, e.g. pixel-wise regression.
In this work, we focus not only on scaling the volume of the pretraining data, but also on scaling the number of parameters and the architecture, studying the transition from CNN-based models to Vision Transformer (ViT) backbones combined with UPerNet decoder. The main contributions of this paper are:
  • We scale the PhilEO pretraining from the PhilEO Globe 0.5TB dataset to the full MajorTOM 23TB dataset, as well as to the specialized FastTOM 2TB subset (i.e. without oceans and ice), with different number of model parameters and architectures [^1] We release the code of our model in: http://github.com/ESA-PhiLab/PhilEO-MajorTOM.. We demonstrate that larger pretraining data reduces the Root Mean Squared Error (RMSE), resulting in a percentage improvement $12.0\%$ for road density regression for the $n$-shot $n=500$.
  • In addition to dataset scaling, we study the scaling of the model parameters and the architecture, comparing CNN- and ViT-based models. We evaluate the effectiveness of the scaling-up on the PhilEO Bench, covering road and building density estimation and land cover mapping.
  • We develop and compare different architectures for the EO Foundation Models in order to select which one to scale-up. For road density regression, we show that the PhilEO ViT UPerNet outperforms the PhilEO ViT with a convolutional decoder in RMSE, with a percentage improvement of $34.2\%$ for the $n$-shot $n=100$. On the same task, the PhilEO Geo-Aware U-Net model achieves the same performance as the PhilEO ViT UPerNet.
The rest of the paper is organized as follows. Section II provides background on the PhilEO Geospatial Foundation Model (GFM) and the PhilEO Bench evaluation, as well as the datasets MajorTOM and FastTOM. Section III describes the proposed PhilEO MajorTOM 23TB and FastTOM 2TB models. Section IV presents the related work. We evaluate our models on the PhilEO Bench in Sec. V for road and building density estimation and land cover mapping. We present the results of the PhilEO MajorTOM 23TB and FastTOM 2TB models, the PhilEO 200M FastTOM, the PhilEO Geo-Aware U-Net, and the PhilEO ViT UPerNet 100M FastTOM.
II Background
PhilEO GFM. The PhilEO model [3] employs a combination of masked reconstruction and geo-location estimation as pretext tasks to train on the PhilEO Globe 0.5TB S2L2A dataset, which contains only land, excluding oceans and ice. PhilEO Globe also includes four time steps: $4$ seasons from the same year. The architecture is a modified U-Net CNN model.
Figure 1: High-level overview of ViT-UPerNet. The dimensions of the feature map $C_{4}$ are maintained. The other feature maps are upsampled, respectively downsampled. The resulting feature map is used for downstream processing.
PhilEO Bench. To enable fair evaluation and comparison of GFMs, [3] introduced the PhilEO Bench which includes a global labeled Sentinel-2 dataset for the standardized downstream tasks: road density estimation, building density regression, and land cover mapping. The bench also incorporates several existing GFMs, such as [6] and [7], allowing for consistent comparisons across models.
MajorTOM and FastTOM datasets. The MajorTOM dataset [15] comprises approximately 60TB of unlabeled global data, covering most of the Earth, including oceans and ice. In this work, we focus on the MajorTOM Core-S2L2A subset, 23TB. FastTOM is a 2TB specialized subset of MajorTOM Core-S2L2A, containing only land and excluding oceans and ice, making it more task-specific for terrestrial downstream tasks.
III Proposed Model
III-A PhilEO MajorTOM 23TB and PhilEO FastTOM 2TB
We extend our previously proposed PhilEO model by pretraining it on the MajorTOM Core-S2L2A dataset [15], which contains 23TB data covering the vast majority of the Earth’s surface, including land, oceans, and ice. To the best of the authors’ knowledge, this is the first instance where the full 23TB MajorTOM dataset is used to pretrain a GFM. The PhilEO MajorTOM model was trained using the Leonardo Davinci-1 Supercomputer, using between four and eight compute nodes, each equipped with four NVIDIA A100 GPUs (40GB VRAM), resulting in a total of 16-32 GPUs. This HPC configuration provided approximately 2.5–5 petaFLOPS of compute performance. Without such infrastructure, training on a single GPU would have taken several months of continuous computation, highlighting the critical role of HPC resources in enabling this work. To handle the scale of the dataset and model, we used PyTorch’s Distributed Data Parallel (DDP) to accelerate training. For larger ViT-based models still ongoing training on MajorTOM, we are transitioning to Fully Sharded Data Parallel (FSDP) to improve memory efficiency and support larger model scales. In addition, mixed-precision training (BF16/BF32) was used for improved throughput and reduced memory usage. Alongside the MajorTOM pretraining, we also trained a model variant on the FastTOM 2TB subset, which contains only land. FastTOM serves as a proxy for MajorTOM: strong performance on FastTOM is expected to translate into good performance when scaling to the full MajorTOM 23TB dataset. This strategy allows for efficient experimentation and architecture selection before undertaking full-scale pretraining.
III-B PhilEO ViT UPerNet
PhilEO Bench [3] standardizes model comparison by using a common convolutional decoder based on a U-Net-like architecture [8]. In this work, we introduce an alternative decoder strategy by implementing the UPerNet decoder [10], shown in Fig. 1, in the PhilEO Bench, in order to compare models with a ViT backbone. The UPerNet design [10] draws inspiration from human visual perception, combining hierarchical feature extraction through a Feature Pyramid Network (FPN) [21] with global context aggregation via a Pyramid Pooling Module (PPM). The FPN aggregates features across multiple scales through a top-down pathway with lateral connections, while the PPM captures context at various spatial resolutions. ViT-UPerNet architectures have been shown to achieve state-of-the-art performance in various computer vision tasks [11]-[14], including segmentation and classification. For models using a ViT backbone, we extract intermediate feature maps, $\{C_{2},C_{3},C_{4},C_{5}\}$ in Fig. 1, across different layers. These feature maps are then appropriately resized (upsampled or downsampled) to align spatial dimensions. The highest-level feature map $C_{5}$ is processed by the PPM, which applies pooling operations at multiple scales to capture contextual information at different resolutions. The outputs of the PPM and the FPN are then fused through a series of 1×1 convolutions and element-wise additions, producing the final set of multi-scale feature maps, $\{P_{2},P_{3},P_{4},P_{5}\}$ in Fig. 1. These aggregated features are then used for downstream pixel-wise prediction tasks, including regression and segmentation.
TABLE I: Comparison of the PhilEO MajorTOM model to other GFMs with respect to the features of the pretraining dataset.
Model Dataset Data size Data features
Prithvi-EO-2.0 Global HLS 4-5TB Land [23]
HydroFM Hydro 1TB Ocean [25]
SatMAE++ fMoW 3.5TB Land [26]
CROMA SSL4EO-S12 1.5TB Land [27]
Copernicus FM C.-Pretrain 10TB Land [28]
TerraMind [29] TerraMesh 14TB Land, ocean, ice
PhilEO Globe [3] Ph. Globe 0.5TB Land
PhilEO MajorTOM MajorTOM 23TB Land, ocean, ice
HLS = Harmonised Landsat S-2; fMoW = Functional Map of the World. Here, the MajorTOM dataset includes oceans and ice, while FastTOM and PhilEO Globe do not.
IV Related Work
Recent studies on GFMs, summarized in Table I, highlight the increasing potential of learning robust feature representations from large-scale Earth Observation (EO) data, later used for fine-tuning on various downstream tasks. However, most existing studies have been limited by relatively modest dataset sizes and have not explored training GFMs at the scale of more than 20TB. This motivates our work, where we develop and pretrain the PhilEO MajorTOM model on the full 23TB MajorTOM Core-S2L2A dataset. In addition, few-shot learning has gained attention within the EO community, as it enables models to generalize effectively from limited labeled data. Prior work [17, 18] demonstrated the advantages of few-shot learning for land cover classification, showing that models can maintain strong performance even with a small number of examples.
The synergy between Foundation Models and few-shot learning was further explored in [19], showcasing improved segmentation performance through combined training strategies. Additionally, frameworks such as [20] have proposed systematic approaches to low- and few-shot adaptation in the context of Foundation Models.
In addition, to this day, U-Net models are still considered the default benchmark for semantic segmentation, often outperforming new architectures. [16] benchmarked the performance of Swin-UPerNet against a U-Net on semantic segmentation use cases. It was shown that the Swin-UPerNet model is a good competitor for the U-Net.
Figure 2: Road density estimation: Evaluation in RMSE at different $n$-shots, where the best performing model is PhilEO Geo-Aware U-Net. The PhilEO ViT UPerNet black line is below the Geo-Aware U-Net red line. PhilEO ViT UPerNet outperforms ViT CNN and ViT CNN Grouped Channels (GC) [3].
Figure 3: Building density regression: Evaluation in RMSE at various $n$-shots. The PhilEO ViT UPerNet black line is below the Geo-Aware U-Net red line.
V Evaluation and Results
We evaluate the proposed models, comparing different architectures and analyzing the impact of scaling-up pretraining. Our evaluation focuses on the three PhilEO Bench downstream tasks: road density estimation, building density regression, and land cover mapping. Before training the models directly on MajorTOM, we first train them on the 0.5TB PhilEO Global dataset in order to select the best architecture to scale-up.
V-A PhilEO ViT UPerNet pretrained on PhilEO Globe 0.5TB
As shown in Fig. 2, the PhilEO ViT UPerNet model consistently outperforms the PhilEO ViT model with a convolutional decoder across all $n$-shot settings for road density estimation. Notably, the PhilEO ViT UPerNet achieves comparable performance to the PhilEO Geo-Aware U-Net model [3], highlighting the effectiveness of the UPerNet decoder for pixel-wise regression tasks.
In Fig. 3, the PhilEO ViT UPerNet demonstrates improved RMSE performance for building density estimation compared to the convolutional decoder ViT baseline. For instance, in Fig. 3, at $n=50$, the PhilEO ViT UPerNet achieves a RMSE of $0.08944$, compared to $0.2$ for the convolutional decoder ViT model, corresponding to an improvement of $55.28$%. At $n=100$, the ViT UPerNet model further improves to a RMSE of $0.08367$, versus $0.1$ for the convolutional decoder ViT baseline, leading to a $16.33$% relative gain.
Figure 4: Land cover mapping: $n$-shot evaluation of models in accuracy.
In Fig. 4, the PhilEO ViT UPerNet also outperforms the ViT CNN model across all $n$-shot experiments for land cover mapping, in accuracy. Moreover, both the PhilEO ViT UPerNet and the Geo-Aware U-Net outperform a fully supervised U-Net model (without pretraining), particularly in low-shot regimes. Overall, the PhilEO Geo-Aware U-Net achieves slightly better performance on average than the PhilEO ViT UPerNet across the three downstream tasks, as shown in Figs. 2-4.
V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
According to the results in Section V-A, we scale-up the best-performing architecture, the PhilEO Geo-Aware U-Net, and we pretrain it on the full 23TB MajorTOM dataset, as well as on the FastTOM 2TB specialized subset. We study the scaling-up to 23TB (and 2TB) data, as well as the number of parameters 200M (versus 44M), and the architecture, i.e. ViT UPerNet (versus Geo-Aware U-Net [3]).
Road density estimation. Fig. 5 shows the RMSE performance for road density estimation across different $n$-shot settings. For all $n$-shots, the scaled-up PhilEO MajorTOM 23TB model outperforms the Geo-Aware U-Net model pretrained on the PhilEO Globe 0.5TB dataset. At $n=100$, the PhilEO MajorTOM 23TB model achieves a RMSE of $0.06724$, compared to $0.07070$ for the model pretrained on the PhilEO Globe 0.5TB dataset. This corresponds to an improvement of $4.89$%. For $n=50$ and $100$, the PhilEO 44M MajorTOM model outperforms the FastTOM and PhilEO Globe pretrained baselines, demonstrating the benefits of large-scale pretraining even when fine-tuning with limited labeled data.
Additionally, we develop and evaluate two larger models, as shown in Fig. 5:
  • PhilEO 200M FastTOM: a 200M-parameter Geo-Aware U-Net variant pretrained on FastTOM,
  • ViT 100M FastTOM: a Vision Transformer model (approximately 100M parameters, with depth $32$, number of heads $16$, and embedding dimension $512$) pretrained on FastTOM with a UPerNet decoder utilizing intermediate layers $5$, $15$, $23$, and $31$.
Both larger models show strong performance, with PhilEO 200M FastTOM outperforming the others at several $n$-shots.
Figure 5: Road density estimation in RMSE, comparing the models: PhilEO Globe 44M, PhilEO MajorTOM and FastTOM, PhilEO 200M FastTOM, ViT UPerNet 100M FastTOM and PhilEO Globe ViT UPerNet 300M. The models PhilEO MajorTOM and FastTOM outperform at most $n$-shots PhilEO Globe.
Figure 6: Building density in RMSE for PhilEO Globe, PhilEO MajorTOM and FastTOM, PhilEO 200M FastTOM, ViT FastTOM and PhilEO Globe ViT.
Building density regression. As shown in Fig. 6, similar trends are observed for building density estimation. At $n=100$ and $500$, both the PhilEO 44M MajorTOM and FastTOM models outperform in RMSE the PhilEO Globe 44M baseline. In addition, the PhilEO 200M FastTOM model again performs best across most $n$-shot settings.
Figure 7: Land cover in accuracy for the different models, at various $n$-shots.
Land cover mapping. Fig. 7 presents accuracy results for land cover semantic segmentation. Across all $n$-shot settings, the PhilEO Geo-Aware U-Net, pretrained on the smaller PhilEO Globe 0.5TB dataset, consistently outperforms models pretrained on the PhilEO MajorTOM and FastTOM datasets. A likely explanation for the superiority of PhilEO Globe in land cover is that it includes seasonal variations, i.e. $4$ seasons, $3$ months separation over $1$ year, $4$ time steps. For land cover, seasonality is beneficial, i.e. seasonality is more correlated with land cover than with roads or buildings: PhilEO Globe includes $4$ time steps (i.e. $1$ year), rather than MajorTOM, one random time step [15]. Also, it is important to note that PhilEO Globe and FastTOM contain only land, while MajorTOM includes a broader variety of scenes, such as land, oceans, ice, forests, and deserts. Since our downstream tasks focus solely on land, the models pretrained on PhilEO Globe and FastTOM benefit from prior knowledge that aligns with the evaluation setting, while MajorTOM-pretrained models do not.
To better understand the impact of specialized versus general-purpose pretraining, we compare the performance of models trained on specialized land-only datasets (PhilEO Globe 0.5TB and FastTOM 2TB) with the performance of models trained on a large general-purpose dataset (MajorTOM 23TB) which includes oceans and ice. Given that all models share the same Geo-Aware U-Net architecture (approximately 44M parameters) and that the smaller datasets use a land mask to exclude oceans and ice, we expect models trained on PhilEO Globe and FastTOM to perform better in downstream land-focused tasks. Indeed, in most cases, this expectation holds. However, for the road density estimation task under low-shot (few-shot) learning settings, the model pretrained on MajorTOM surprisingly outperforms the specialized models in terms of RMSE. This suggests that despite the broader and noisier pretraining data, large-scale general pretraining can be advantageous under certain conditions.
This study allows us to analyze two key factors:
  • Specialization effect: Does pretraining on land-only images give an advantage when downstream tasks are also land-focused?
  • Scaling effect: How does increasing the pretraining dataset size from 0.5TB and 2TB (specialized) to 23TB (general-purpose) impact downstream performance?
We study how GFMs scale and explore how the PhilEO models behave, their scaling behavior. Pretraining on MajorTOM represents a truly general-purpose approach, aligning with the concept of a GFM for Earth observation, rather than a Land-only model. GFMs pretrained on general-purpose satellite data, and not only on land, are Earth Observation models and not only Land Observation models (or Ocean Observation models). In contrast, models like SeCo [24] focus on land observations, while others, like HydroFM [25] (Table I), specialize in oceans. Our results show that although specialized pretraining offers an initial advantage, scaling-up the pretraining on diverse general-purpose datasets, when scaling-up is performed appropriately, can eventually lead to GFMs that outperform specialized models, even in tasks where prior domain-specific knowledge initially seemed crucial.
Figure 8: Land cover mapping: PhilEO MajorTOM model precision per class.
Figure 9: Land cover mapping: PhilEO FastTOM model precision per class.
Per-class results for land cover mapping. We evaluate and examine the precision metrics for each class across different $n$-shots settings in Fig. 8, focusing on the PhilEO 44M MajorTOM 23TB model’s performance on the PhilEO Bench land cover semantic segmentation task, which includes $11$ classes based on the ESA WorldCover classification scheme. Similarly, Fig. 9 presents the class-wise precision metrics for the PhilEO 44M FastTOM 2TB model on the same land cover semantic segmentation task. The results from both figures demonstrate that precision generally improves across most classes as the number of $n$-shot examples increases.
VI Conclusion
In this work, we studied scaling-up the pretraining data for GFMs. By scaling the pretraining dataset from 0.5TB (PhilEO Globe) to 23TB (MajorTOM), we achieved improvements in road and building density regression when labeled data was scarce. For land cover mapping, models pretrained on the specialized land-only dataset PhilEO Globe performed better than the MajorTOM-pretrained model, due to the prior exposure to both seasonality and land-specific imagery. Our study confirms that pretraining on large, diverse datasets—combined with appropriate architectural choices—enables the development of more robust and generalizable GFMs. The results underline the critical role of high-performance computing infrastructure and optimized training strategies in scaling-up EO AI systems, paving the way for future work on even larger models and more complex downstream applications.